OpenStudy (anonymous):

The position of a particle is given by: s(t) = -e^(-t) sin(t) a)find a formula for velocity of a particle b) is there a time t when the particle changes direction? if so give a particular value of t where this happens. c) what is the limit of the velocity of the particle as t tends to infinity?

OpenStudy (anonymous):

velocity means derivative take the derivative to get the answer to A

OpenStudy (anonymous):

\[v(t) = -e^{-t}\sin(t) +e ^{-t}\cos(t)\]

OpenStudy (anonymous):

ok i believe you set it equal to zero for part B

OpenStudy (anonymous):

\[0 =(-\sin(t)+\cos(t)) e ^{-t}\]

OpenStudy (anonymous):

let me check your derivative before we go further

OpenStudy (anonymous):

check it again i htink it is wrong brb

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

not that it will make any difference when we set it equal to zero and solve, but i think the derivaitve is \[e^{-t}(\sin(t)-\cos(t))\]

OpenStudy (anonymous):

in any case you want \[\sin(t)-\cos(t)=0\\ \sin(t)=\cos(t)\\ t=\frac{\pi}{4}\] etc

OpenStudy (anonymous):

oh okay that makes sense

OpenStudy (anonymous):

but then how do i get part c? do i just plug in really big numbers?